# Orthogonalized Policy Optimization Decoupling Sampling Geometry from Optimization Geometry in RLHF

### v3

@@ 5.1 Formal Decoupling of Geometry Axes
Direct differentiation yields a constant Hessian $\mu I$, which is independent of $\alpha$. ∎
Remark 5.3 (Scope of Orthogonality).
The structural orthogonality in Theorem 5.2 holds exactly in the ratio function space $L^{2}(\pi_{\mathrm{ref}})$. In parameter space $\theta$, interaction may arise through policy parametrization, since $\nabla_{\theta}v=\nabla_{\theta}(\exp(\Delta)-1)$ depends on the training trajectory influenced by $\alpha$. However, this coupling is a property of the parametrization, not the intrinsic objective curvature: the Hessian $\nabla^{2}_{v}\mathcal{L}=\mu I$ remains structurally independent of $\alpha$ regardless of the parameter mapping.
A natural concern is whether the stop-gradient operator in $\omega_{\alpha}$ (Remark 4.3) introduces bias. The following lemma shows it does not, at the anchor point:
Lemma 5.4 (Exactness of Surrogate Gradient at Anchor Point).
Let $\mathcal{L}_{\mathrm{true}}(\theta)$ be the exact objective without stop-gradient on $\omega_{\alpha}$, and $\mathcal{L}_{\mathrm{OPO}}(\theta)$ the surrogate with $\operatorname{sg}$. Under on-policy anchoring $\pi_{\mathrm{ref}}=\pi_{\mathrm{old}}$, at the start of each iteration ($\theta=\theta_{\mathrm{old}}$):
$\left.\nabla_{\theta}\mathcal{L}_{\mathrm{OPO}}(\theta)\right|_{\theta=\theta_{\mathrm{old}}}=\left.\nabla_{\theta}\mathcal{L}_{\mathrm{true}}(\theta)\right|_{\theta=\theta_{\mathrm{old}}}$ (17)
Furthermore, the gradient error grows only as $\mathcal{O}(\|\pi_{\theta}-\pi_{\mathrm{old}}\|)$.
Proof.
By the product rule, the true function-space gradient contains a cross-term $(\nabla_{v}\omega_{\alpha})\,v$ that the surrogate omits. At the anchor point $\pi_{\theta}=\pi_{\mathrm{old}}$, the ratio offset is exactly $v=\pi_{\theta}/\pi_{\mathrm{old}}-1=0$, so this cross-term vanishes. For $\theta\neq\theta_{\mathrm{old}}$, the error is bounded by $\|(\nabla_{v}\omega_{\alpha})\,v\|=\mathcal{O}(\|v\|)=\mathcal{O}(\|\pi_{\theta}-\pi_{\mathrm{old}}\|)$, which is small within the trust region. ∎
This confirms that stop-gradient is not an engineering hack but a first-order exact approximation: the surrogate and true objectives agree to first order at every on-policy anchor. Under sufficiently small trust-region updates (i.e., $\|v\|$ bounded by the $\chi^{2}$ penalty), the cumulative bias across iterations remains second-order small, since each anchor reset re-zeroes the first-order error term.
Corollary 5.5 (Global Contraction and Linear Convergence).
Under the assumption that the policy class can represent $v^{*}$ and that local linearization of the map $\theta\mapsto v_{\theta}$ holds near the optimum, gradient descent in $v$-space on the OPO objective follows a linear dynamical system:
$v_{k+1}-v^{*}=(1-\eta\mu)(v_{k}-v^{*})$ (18)
where $v^{*}=\omega_{\alpha}/\mu$. For step size $0<\eta<2/\mu$, this system exhibits global contraction to the unique equilibrium. The contraction rate $|1-\eta\mu|$ depends solely on the optimization stiffness $\mu$, confirming that convergence speed is structurally decoupled from the target distribution sharpness $\alpha$.
We use the term “orthogonalized” to emphasize this control decoupling.
Remark 5.6 (Contrast with KL-Based Methods).
In KL-regularized objectives (e.g., DPO), the effective loss locally resembles a logistic function $\mathcal{L}\approx-\log\sigma(\text{margin})$. The local curvature scales as $\beta^{2}\sigma(m)(1-\sigma(m))$, where $m$ is the logit margin. This curvature depends on both the temperature $\beta$ and the current data margin $m$. Thus, changing $\beta$ simultaneously alters both the gradient magnitude and the local stability profile, entangling the two design axes. OPO’s constant curvature $\mu$ avoids this data-dependent instability.
5.2 Comparison with Parametric L2 Regularization
A natural question is whether OPO differs fundamentally from standard Policy Gradient with L2 parameter regularization (weight decay).
$\mathcal{L}_{\text{L2-PG}}=-\mathbb{E}[A\log\pi_{\theta}]+\frac{\lambda}{2}\|\theta-\theta_{0}\|^{2}$ (19)
While both methods induce a quadratic penalty, they operate in fundamentally different spaces:
  1. 1.
    Functional vs. Parametric Prior: OPO’s penalty $\frac{\mu}{2}\mathbb{E}[v^{2}]$ acts in the function space of probability ratios. It penalizes deviations only where the policy assigns probability mass. L2 parameter regularization acts in the weight space. Two different policies can have the same L2 distance, and small weight changes can cause large output deviations (and vice-versa).
  2. 2.
    Adaptive Constraint: OPO’s regularization provides a “linear restoring force” directly on the probability ratio. If the model is confident (large $v$), the $\chi^{2}$ penalty generates a strong, specific gradient signal to pull that specific output back. Parametric L2 is blind to the output; it uniformly decays all weights regardless of their contribution to the current output confidence.
  3. 3.
    Learning Dynamics: Empirically, Functional $\chi^{2}$ regularization allows the model to shift mass aggressively where advantages are high, while maintaining stability. Weight decay suppresses all weight magnitudes, which can inadvertently hinder the learning of sharp, high-confidence reasoning chains required for math tasks.
5.3 Quantitative Gradient Analysis
We provide a quantitative comparison of the gradient signal in high-confidence regimes, proving the non-saturation property.
Gradient Saturation in DPO.
For logistic-based losses (DPO), the gradient magnitude with respect to the margin $m$ is $|\nabla_{m}\mathcal{L}|=|\beta\sigma(m)(1-\sigma(m))|$. Since $\sigma(1-\sigma)\leq 1/4$:
$|\nabla_{\text{DPO}}|\leq\frac{\beta}{4}$ (20)
Crucially, as the margin $m\to\infty$, $|\nabla_{\text{DPO}}|\to 0$ exponentially.
Linear Gradient Dynamics in OPO.
For OPO, the gradient is $\nabla_{v}\mathcal{L}=\mu v-\omega_{\alpha}$. Let $v^{*}=\omega_{\alpha}/\mu$. The gradient magnitude is exactly proportional to the distance from equilibrium: $|\nabla_{v}\mathcal{L}|=\mu|v-v^{*}|$. Therefore, for any non-equilibrium state ($|v-v^{*}|\geq\delta$):
$|\nabla_{\text{OPO}}|\geq\mu\delta$ (21)
This guarantees a linear driving force that does not vanish arbitrarily. The practical consequence is stark: in mathematical reasoning, models frequently assign near-1.0 probability to correct intermediate steps. KL-based methods penalize these high-confidence correct outputs by exponentially shrinking the learning signal, whereas OPO maintains a constant reward-driven pull via its linear dynamics—allowing the model to continue refining even its most confident predictions.
@@ 5.4 Algorithm and Flowchart
We provide a visual overview of the OPO process in Figure 1 and the detailed procedure in Algorithm 1.
@@
Figure 1: Flowchart of the Orthogonalized Policy Optimization (OPO) framework. The “OPO Decomposition” block (green) embodies the two-geometry structure: the $\alpha$-divergence projection (step 2) governs sampling, while the $\chi^{2}$ penalty (step 3) governs optimization. These two axes are structurally orthogonal.
@@
5:  Advantage: $A(y)\leftarrow\mathrm{Normalize}(R(x,y))$ $\triangleright$ Group-normalized
@@
8:  Loss: $\mathcal{L}\leftarrow\frac{1}{N}\sum_{y}\big[-\omega_{\alpha}(y)\,\Delta_{\theta}(y)+\frac{\mu}{2}\,\Delta_{\theta}(y)^{2}\big]$ $\triangleright$ $y$: complete response sequence
@@ 6.1 Experimental Setup
  • GSPO (Group Sentence-level Policy Optimization): A variant of GRPO that applies advantage normalization at the sentence/step level rather than the token level, aiming for more granular credit assignment.
  • DAPO (Dynamic Advantage Policy Optimization): An advanced baseline incorporating four key mechanisms:
@@
      Dynamic Sampling: dynamically filters "all-correct" or "all-wrong" samples that contribute zero advantage, focusing training on informative samples.
@@ 6.2 Results
The experimental results confirm that OPO’s orthogonalized design—combining $\alpha$-weighted sampling with $\chi^{2}$ optimization geometry—prevents gradient saturation and enables continued learning in high-confidence regimes.

### v4

@@ 4.4 Closed-Form Solution (Linear Response Law)
Introduce a Lagrange multiplier $\lambda$ for $\mathbb{E}_{\pi_{k}}[v]=0$. Stationarity gives $0=-\omega_{\alpha}(y)+\mu v(y)+\lambda$. Taking $\mathbb{E}_{\pi_{k}}[\cdot]$ and using $\mathbb{E}_{\pi_{k}}[v]=0$ yields $\lambda=\mathbb{E}_{\pi_{k}}[\omega_{\alpha}]=1$. Substituting back gives $v^{*}=(\omega_{\alpha}-1)/\mu$. ∎
Policy update.
Since $\pi=\pi_{k}(1+v)$,
$\frac{\pi_{k+1}(y)}{\pi_{k}(y)}=1+\frac{1}{\mu}\Big(\omega_{\alpha}(y)-1\Big).$ (12)
Non-saturating linear dynamics.
The function-space gradient is $\nabla_{v}\mathcal{L}_{\mathrm{OPO}}=-\omega_{\alpha}+\mu v$, which is linear in $v$ with constant curvature $\nabla_{v}^{2}\mathcal{L}=\mu I$. Unlike sigmoid/logistic objectives, this does not exhibit exponential saturation.
Figure 1: Work–dissipation physics picture of OPO. The anchor policy $\pi_{k}$ sits at the minimum of a quadratic fluctuation energy ($\chi^{2}$ geometry). The escort field $\rho_{\alpha}$ induces the driving weight $\omega_{\alpha}=\rho_{\alpha}/\pi_{k}$. The perturbed policy $\pi$ corresponds to a ratio fluctuation $v=\pi/\pi_{k}-1$. The conservation law $\mathbb{E}_{\pi_{k}}[v]=0$ forces an orthogonal projection, yielding the centered response $v^{*}=(\omega_{\alpha}-1)/\mu$. In statistical mechanics [18], this is the fluctuation–dissipation theorem: response equals susceptibility ($1/\mu$) times centered driving force.
5 Mirror Descent Interpretation
We show that the work–dissipation principle is equivalently a mirror descent step with a Euclidean mirror map.
5.1 Euclidean Mirror Map and $\chi^{2}$ Bregman Divergence
Define the mirror map $\Psi(v):=\frac{1}{2}\|v\|_{L^{2}(\pi_{k})}^{2}$. The induced Bregman divergence is
$D_{\Psi}(\pi\|\pi_{k})=\frac{1}{2}\mathbb{E}_{\pi_{k}}\!\left[\left(\frac{\pi}{\pi_{k}}-1\right)^{2}\right],$ (13)
which is the Pearson $\chi^{2}$ divergence in ratio coordinates.
5.2 Orthogonal Mirror Step
A mirror/proximal step driven by $\omega_{\alpha}$ is
$v^{*}=\operatorname*{arg\,min}_{v\in\mathcal{H}_{0}}\left\{-\langle\omega_{\alpha},v\rangle+\frac{\mu}{2}\|v\|^{2}\right\},$ (14)
which is identical to the work–dissipation formulation (9). Thus mirror descent and the free-energy principle define the same update. $\alpha$ affects only the drive $\omega_{\alpha}$; $\mu$ affects only the proximal curvature.
6 Hilbert Projection Interpretation
We re-derive the same objective from the Hilbert space projection theorem, requiring no mirror maps.
6.1 Hilbert Space and Metric Modulation
Let $\mathcal{H}=L^{2}(\pi_{k})$ with inner product $\langle f,g\rangle=\mathbb{E}_{\pi_{k}}[fg]$. Define the multiplication operator (diagonal preconditioner)
$(\mathcal{M}_{\alpha}f)(y):=\omega_{\alpha}(y)\,f(y).$ (15)
Lemma 6.1 (Boundedness, Self-Adjointness, Positivity).
If $\omega_{\alpha}\in L^{\infty}(\pi_{k})$ and $\omega_{\alpha}>0$ a.s., then $\mathcal{M}_{\alpha}$ is bounded, self-adjoint, and positive. The induced inner product $\langle f,g\rangle_{\alpha}:=\mathbb{E}_{\rho_{\alpha}}[fg]$ is the escort-modulated geometry.
Thus $\alpha$-escort is a smooth metric modulation: it preserves support and reweights directions in function space through a new local inner product, in contrast to hard masking which truncates support.
6.2 Projection Theorem Gives the Same Closed Form
Rewrite Eq. (9) by completing the square:
$\langle\omega_{\alpha},v\rangle-\frac{\mu}{2}\|v\|^{2}=-\frac{\mu}{2}\left\|v-\frac{1}{\mu}\omega_{\alpha}\right\|^{2}+\frac{1}{2\mu}\|\omega_{\alpha}\|^{2}.$ (16)
Hence
$v^{*}=\operatorname*{arg\,min}_{v\in\mathcal{H}_{0}}\left\|v-\frac{1}{\mu}\omega_{\alpha}\right\|^{2}.$ (17)
By the Hilbert projection theorem,
$v^{*}=\frac{1}{\mu}\mathcal{P}_{\mathcal{H}_{0}}(\omega_{\alpha})=\frac{1}{\mu}\big(\omega_{\alpha}-\langle\omega_{\alpha},1\rangle\,1\big)=\frac{1}{\mu}(\omega_{\alpha}-1),$ (18)
recovering Theorem 4.3. The subtraction of $1$ is the unique orthogonal correction enforced by conservation.
Remark 6.2 ($z$-Score Normalization as Conservation-Law Projection).
The centering $\omega_{\alpha}\mapsto\omega_{\alpha}-\mathbb{E}[\omega_{\alpha}]$ is exactly $\mathcal{P}_{\mathcal{H}_{0}}$—the mean-subtraction step of $z$-score normalization. When $\omega_{\alpha}(y)=A(y)$ ($\alpha\to 1$), the solution $v^{*}=(A-\mathbb{E}[A])/\mu$ is the $z$-scored advantage scaled by susceptibility. Thus, advantage $z$-scoring is not a variance-reduction heuristic: it is the unique orthogonal projection enforcing probability conservation in $L^{2}(\pi_{k})$.
Remark 6.3 (Contrast with Fisher–Rao (KL) Geometry).
In KL-based methods, the natural metric is the Fisher–Rao metric $g_{ij}^{\mathrm{FR}}=\mathbb{E}_{\pi}[\partial_{i}\log\pi\cdot\partial_{j}\log\pi]$, whose curvature depends on $\pi$. In contrast, $\chi^{2}$ geometry operates in $L^{2}(\pi_{k})$ with constant inner product:
KL / Fisher–Rao $\chi^{2}$ / $L^{2}(\pi_{k})$
Inner product Fisher metric (policy-dependent) $L^{2}(\pi_{k})$ (fixed at anchor)
Hessian $\nabla^{2}\propto\beta^{2}\sigma(1{-}\sigma)$ $\nabla^{2}=\mu I$ (constant)
Perturbation theory Nonlinear (exponential) Linear (first-order exact)
7 Implementable Surrogate and Algorithm
7.1 Log-Ratio Surrogate for LLMs
LLMs output log-probabilities. Let $\Delta_{\theta}(y):=\log\pi_{\theta}(y)-\log\pi_{k}(y)$. Since $v=e^{\Delta}-1\approx\Delta$ for small trust regions, the practical surrogate is
$\boxed{\mathcal{L}_{\mathrm{OPO}}^{\log}(\theta)=-\mathbb{E}_{\pi_{k}}\big[\omega_{\alpha}(y)\,\Delta_{\theta}(y)\big]+\frac{\mu}{2}\mathbb{E}_{\pi_{k}}\big[\Delta_{\theta}(y)^{2}\big].}$ (19)
Anchoring $\pi_{k}=\pi_{\mathrm{old}}$ resets $\Delta$ near zero each iteration, supporting the local approximation.
@@ 7.2 Algorithm and Flowchart
Figure 2: Flowchart of the OPO framework. The “OPO Decomposition” block (green) embodies the work–dissipation structure: the $\alpha$-escort (step 2) shapes the external work, while the $\chi^{2}$ penalty (step 3) shapes the intrinsic dissipation. The two axes are structurally orthogonal.
@@
5:  Advantage: $A(y)\leftarrow\mathrm{zscore}(R(x,y))$ $\triangleright$ Group $z$-score (Remark 6.2)
@@
8:  Loss: $\mathcal{L}\leftarrow\frac{1}{N}\sum_{y}\big[-\omega_{\alpha}(y)\,\Delta_{\theta}(y)+\frac{\mu}{2}\,\Delta_{\theta}(y)^{2}\big]$
@@
8 Theoretical Analysis
8.1 Constant Curvature and Orthogonal Control
Theorem 8.1 (Decoupling of Sampling and Optimization Geometry).
Consider the OPO objective in $v$-space. Then:
  1. 1.
    The first-order drive depends on $\alpha$ only through $\omega_{\alpha}$: $\nabla_{v}\mathcal{L}_{\mathrm{OPO}}=-\omega_{\alpha}+\mu v$.
  2. 2.
    The second-order curvature is constant and independent of $\alpha$: $\nabla_{v}^{2}\mathcal{L}_{\mathrm{OPO}}=\mu I$.
Changing $\alpha$ reshapes the target (what to learn) without altering curvature (how to learn). Changing $\mu$ alters curvature without reshaping the target.
8.2 Global Contraction in $v$-Space
Gradient descent $v^{(m+1)}=v^{(m)}-\eta(\mu v^{(m)}-\omega_{\alpha})$ satisfies
$v^{(m+1)}-v^{*}=(1-\eta\mu)\,(v^{(m)}-v^{*}).$ (20)
Corollary 8.2 (Linear Rate).
For $0<\eta<2/\mu$, $\|v^{(m)}-v^{*}\|\leq|1-\eta\mu|^{m}\,\|v^{(0)}-v^{*}\|$. The contraction rate depends only on $\mu$ and is independent of $\alpha$.
8.3 Why KL-Based Objectives Saturate
In KL-regularized preference optimization, local curvature scales like $\sigma(m)(1-\sigma(m))$ and vanishes exponentially as $|m|\to\infty$. By contrast, OPO has constant curvature $\mu I$ and linear restoring force $\|\nabla_{v}\mathcal{L}\|=\|\mu(v-v^{*})\|$, which remains proportional to distance-to-equilibrium.
Figure 3 visualizes this contrast on a binary logit toy model ($z_{\mathrm{ref}}=0$). Panel (a) plots $|\partial\mathcal{L}/\partial z|$ versus the model confidence $p=\sigma(z)$: SFT, DPO, GRPO, DAPO, and GSPO all decay to zero in the high-confidence regime, while OPO’s $\chi^{2}$ gradient grows linearly. Panel (b) zooms into $p\geq 0.5$ on a logarithmic scale, revealing that clipped objectives (GRPO, DAPO, GSPO) suffer a hard cliff to machine-epsilon once the ratio exits the clip window, and DPO saturates exponentially. OPO is the only objective whose gradient magnitude remains $O(1)$ throughout.
Figure 3: Toy gradient saturation experiment (binary logit, $z_{\mathrm{ref}}=0$). (a) Gradient magnitude vs. model confidence: OPO ($\chi^{2}$, green) exhibits linear growth while all other objectives saturate. (b) Log-scale zoom on $p\geq 0.5$: clipped objectives cliff-drop by up to 9 orders of magnitude; OPO maintains non-vanishing gradients throughout.
8.4 Trust-Region Duality
Proposition 8.3 ($\chi^{2}$ Trust-Region Duality).
The constrained problem $\max_{v\in\mathcal{H}_{0}}\langle\omega_{\alpha},v\rangle$ s.t. $\|v\|^{2}\leq\varepsilon$ has Lagrangian relaxation $\max_{v\in\mathcal{H}_{0}}\langle\omega_{\alpha},v\rangle-\frac{\mu}{2}\|v\|^{2}$, with $\mu$ as the multiplier. Moreover, $\|v\|^{2}=\mathbb{E}_{\pi_{k}}[(\pi/\pi_{k}-1)^{2}]$ is the Pearson $\chi^{2}$ divergence.
8.5 Positivity Safeguards
The closed-form linear response $v^{*}(y)=(\omega_{\alpha}(y)-1)/\mu$ is derived in the tangent space of the probability manifold at $\pi_{k}$, and only guarantees first-order stationarity of the proximal objective. It does not enforce $\pi_{k}(y)(1+v^{*}(y))\geq 0$ outside a local neighborhood. When $\omega_{\alpha}(y)\ll 1$ and $\mu$ is small, one may have $v^{*}(y)<-1$, violating positivity.
Proposition 8.4 (Local Positivity Condition).
The update $\pi_{k+1}=\pi_{k}(1+v^{*})$ is a valid probability distribution whenever $\mu\geq\|\omega_{\alpha}\|_{\infty}-1$, since then $v^{*}(y)=(\omega_{\alpha}(y)-1)/\mu\geq-1$ for all $y$.
In practice, positivity is maintained by one or more of: (i) choosing $\mu$ large enough relative to the escort weight range; (ii) clipping $v^{*}$ to $1+\epsilon,\,+\infty)$; (iii) projecting onto the probability simplex after the update. This is consistent with standard trust-region Taylor approximations in policy optimization and does not alter local convergence properties (Corollary 8.2).
@@ 9.1 Experimental Setup
  • GSPO (Group Sequence Policy Optimization) [21]: Applies sequence-level importance ratios rather than token-level ratios, aiming for more granular credit assignment at the trajectory level.
  • DAPO (Decoupled Clip and Dynamic sAmpling Policy Optimization) [20]: An advanced baseline incorporating four key mechanisms:
@@
      Dynamic Sampling: dynamically filters “all-correct” or “all-wrong” samples that contribute zero advantage, focusing training on informative samples.
@@ 9.2 Results
Gradient Saturation Analysis.
Figure 6 provides direct evidence for OPO’s non-saturating property. Panel (a) plots gradient norms against mean reward (a confidence proxy). As the model becomes more confident (higher reward), GRPO and GSPO gradients exhibit a decaying trend, and DAPO gradients remain near zero. In contrast, OPO maintains elevated gradient norms across all confidence levels.
Panel (b) addresses the fact that different methods operate at different advantage scales (e.g., GRPO $z$-scores advantages to $\sigma\!\approx\!2$, whereas OPO uses raw advantages with $\sigma\!\approx\!0.2$). To enable a fair comparison, we plot the gradient response efficiency $\|\nabla\|/\bigl(\tfrac{1}{2}\text{AdvRange}\bigr)$ over training steps, normalizing gradient magnitude by the advantage signal amplitude. OPO sustains an efficiency of ${\sim}1.0$$1.5$ throughout training, approximately $5{\times}$ higher than GRPO (${\sim}0.2$$0.3$) and DAPO (${\sim}0.1$). This confirms the linear response prediction of Theorem 4.3: under $\chi^{2}$ geometry, each unit of advantage signal produces a proportional gradient update, whereas KL-based objectives attenuate the signal through their policy-dependent curvature.
Figure 6: Gradient saturation analysis. (a) Gradient norm vs. mean reward (confidence proxy): OPO maintains non-saturating gradients as confidence grows, while baselines decay. (b) Gradient response efficiency (gradient norm normalized by advantage amplitude) over training: OPO sustains ${\sim}5{\times}$ higher gradient-per-signal than baselines, confirming the constant-curvature prediction of $\chi^{2}$ geometry.
@@
The experimental results confirm that OPO’s orthogonalized design—combining $\alpha$-weighted sampling with $\chi^{2}$ optimization geometry—prevents gradient saturation and enables continued learning in high-confidence regimes. The gradient saturation analysis (Figure 6) provides the most direct evidence: OPO’s gradient magnitude remains proportional to advantage amplitude (linear response), while KL-based baselines exhibit suppressed gradients regardless of the driving signal.
@@ 10 Discussion
Geometry Comparison.
Table 2 summarizes the structural differences among representative methods.
Table 2: Comparison of sampling and optimization geometries across alignment methods.
Method Sampling Geometry Optimization Geometry
PPO / GRPO implicit KL KL curvature (ratio clip)
DPO KL KL curvature (sigmoid)
OPO (Ours) $\alpha$-escort (tunable) $\chi^{2}$ curvature (constant $\mu I$)
Geometric Interpretation of Entropy Preservation.
From a geometric perspective, KL geometry inherits a logarithmic penalty on low-probability outcomes: $\log\pi(y)$ diverges as $\pi(y)\to 0$, strongly suppressing entropy by penalizing any residual mass on unlikely tokens. In contrast, $\chi^{2}$ geometry penalizes deviations quadratically in ratio space: the cost $(\pi/\pi_{k}-1)^{2}$ remains finite even when $\pi(y)\ll\pi_{k}(y)$, which attenuates the shrinkage of low-probability mass and thus helps maintain entropy. This geometric explanation complements the sequence-level vs. token-level distinction discussed in the experiments.
@@
  • Mini-batch anchoring. Within one outer iteration, $\pi_{\mathrm{ref}}$ is frozen at the policy snapshot taken before the first mini-batch gradient step. All subsequent mini-batches in the same iteration share this anchor, ensuring that the log-ratio $\Delta_{\theta}=\log\pi_{\theta}-\log\pi_{\mathrm{ref}}$ is evaluated against a consistent reference. This is analogous to PPO’s “epoch” convention and guarantees that $\Delta_{\theta}$ starts near zero at the beginning of each outer iteration, supporting the local linear approximation $v\approx\Delta$.